OpenStudy (anonymous):

What is the graph of the function f(x)=1/2 x^2

OpenStudy (lanhikari22):

It's the same plot for f(x) = x^2, except, for every y value you get, you divide that by 2. This is called stretching. There's nothing special about graphing it, follow the general rules. f(x) = 1/2 x^2. what's f(2) ? That'd be 1/2 * 2^2 = 1/2 * 4 = 2 g(x) = x^2. what's g(2) ? 2^2 = 4! Here: play with this if it helps.... https://www.desmos.com/calculator/0gzkrjr7mo

OpenStudy (anonymous):

I have no idea how to do any of that. I am so far behind, that im really struggling.

OpenStudy (mathmale):

f(x)=1/2 x^2 would be clearer if written as \[f(x)=\frac{ 1 }{ 2 }x^2\]

OpenStudy (mathmale):

This is an example of "quadratic equations," which you can recognize easily because the highest power of x is 2. This is also classified as a "polynomial," even tho' it has only one term. There are a whole bunch of basic functions that would be worthwhile to learn and review frequently, including their respective graphs. The basic graph of y = x^2 is